# AutoCV: Empowering Reasoning with Automated Process Labeling via Confidence Variation

### v2

@@ 4.2.2 Quantitative Results
We employ three metrics for a thorough evaluation of process calculation hallucination detection:
Precision, the ratio of samples accurately identified as true positives to the total number of positive predictions detected by the OSV model via confidence variations. In our setting, this metric calculates the proportion of samples that have correct final answers but exhibit hallucinatory errors during the reasoning process, relative to the total examples detected by the OSV model.
Recall, which determines the proportion of samples with hallucinatory computational errors that the OSV model successfully identifies through confidence variations;
F1-score, which gauges the verifier’s overall efficacy. In Table 5, we explore how different threshold ($\theta$) values affect the accuracy, recall, and F1-score for process calculation hallucination detection.
Table 5: Process Calculation Hallucination Detection Performance with Varying Threshold ($\theta$) Values.
@@
The results in Table 5 demonstrate that our method using confidence variation effectively detects calculation errors across threshold values from - 0.5 to - 0.9. As the threshold becomes more negative (stricter for labeling errors), the precision increases, indicating higher precision in identifying true errors. However, the recall decreases, meaning fewer actual errors are caught. Importantly, the F1-score, balancing precision and recall, remains relatively stable across thresholds. This demonstrates that our method strikes a good balance between detecting real errors and minimizing incorrect flagging of valid calculations. Overall, our detection method is effective and robust, performing well over a range of thresholds without significantly compromising overall detection quality.
We find that setting $\theta$ = - 0.5 in our detection methods helps maintain a balance between precision and recall, which can ensure a balanced distribution of labeled “incorrect” and “correct” responses.
4.3 Validation and Foundation for AutoCV
By validating these two points, we establish an experimental basis for automating process annotations using our proposed method, AutoCV. Meanwhile, Theorem 1 provides the theoretical groundwork for leveraging the OSV model to estimate the likelihood of arriving at a correct final answer from an intermediate reasoning step onward. These combined theoretical and empirical insights lay a solid foundation for applying AutoCV to construct process-supervision training data for the main experiments that follow.
@@ 5.1 Experimental Setup
Models: We selected three different LLMs of varying sizes, ranging from 7 billion to over 70 billion parameters, to serve as the response generator. Specifically, we used Mistral-Instruct-7B (Mistral-Instruct), Mixtral-8x7B-Instruct-v0.1 (Mixtral-Instruct), and Qwen-72B-Chat (Qwen). Note that all of these are instruction fine-tuning versions.
Datasets: We conduct experiments on five datasets. For mathematical reasoning, we include GSM8K [21], a dataset containing math word problems requiring multi-step reasoning, and MATH [38], composed of high school-level competition problems covering a range of math subjects. For commonsense reasoning, we use HellaSwag [39], a dataset for physically situated commonsense reasoning, Winogrande [40], fill-in-the-blank problems requiring commonsense pronoun resolution and ANLI [41], a dataset for natural language understanding and reasoning.
Evaluation: We follow evaluation metrics in [42] to ensure consistency across all benchmarks. To ensure more reliable pass@$k$ results, we utilized the estimation method described in [43]. We generated $n$ samples per task, where $n$ is greater than $k$, and evaluated the number of correct samples that passed unit tests. We then calculated the unbiased estimator for pass@$k$. For the self-consistency (Self-Cons.) and verifier’s selection results, we randomly choose $k$ out of $n$ samples and conducted separate calculations. The results are reported with an accuracy of ±0.1 at a 95% confidence level. Further details are provided in Section F.2.
@@ 5.2 Enhanced LLMs Reasoning via Process Supervision
We report experimental results on both mathematical and commonsense reasoning tasks across five datasets to showcase the efficacy and scalability of our proposed approach as described in Section 5.2.
Specifically, we calculate $\Delta_{conf}^{t}$ based on the model confidence from OSV and set the threshold $\theta$ = - 0.5 to annotate process-supervision training data autonomously. We further leverage this process-supervision training data to fine-tune the OSV model continually, denoted as OSV + PSV.
For a comprehensive evaluation of our framework, we demonstrate its impact on both mathematics reasoning (GSM8K and MATH datasets) and commonsense reasoning (HellaSwag, Winogrande, and ANLI datasets). We note that Pass@5 represents the upper limit of performance on these benchmarks. We compare the performance of three models: Self-Consistency (Self-Cons.), outcome-supervised verifier (OSV), and process-supervised enhanced verifier (OSV + PSV).
@@
Mistral-7B 69.90 50.03 61.18 61.41 7.7 1.64 5.10 5.30
Mixtral-8$\times$7B 82.30 69.06 74.91 76.04 22.80 10.66 15.2 16.92
Qwen-72B 91.13 81.27 84.91 85.15 56.10 40.10 38.94 39.36
@@
Mistral-7B 76.84 40.30 73.81 74.45 91.16 58.64 79.16 79.98 73.4 45.6 59.8 59.3
Mixtral-8$\times$7B 84.05 73.67 82.83 83.62 79.16 68.75 73.40 73.88 68.4 59.0 62.9 64.0
@@
Mathematics Reasoning: As shown in Table 6, the process-supervised enhanced verifier demonstrates superior performance over the outcome-supervised verifier and Self-Consistency models for all evaluated response generators on GSM8K. For the MATH benchmark, the process-supervised enhanced verifier outperforms the other two approaches for Mistral-Instruct and Mixtral-Instruct, but it is slightly less effective than the Self-Consistency model when applied to Qwen-72b.
Commonsense Reasoning: According to Table 7, OSV + PSV again leads to the best results among the three methods for each response generator tested on HellaSwag. For Winogrande, Mistral-Instruct paired with OSV + PSV achieves the highest performance, whereas, for Mixtral-Instruct and Qwen-72b, the original OSV without process supervision has a marginal advantage. When looking at the results of the ANLI benchmark, OSV + PSV is the highest-performing method for Mistral-Instruct and Mixtral-Instruct. Despite this, for Qwen-72b, the OSV model alone falls slightly behind the integrated OSV + PSV.
In summary, our process-supervision enhanced OSV model (OSV + PSV) consistently matches or improves upon the OSV and Self-Consistency baselines across most benchmarks and response generator models, demonstrating the effectiveness of the automatic process annotation technique in enhancing the verifier’s capabilities for different reasoning tasks.
@@ 6 Analysis
In Section 6.1, we compare our process annotation method, AutoCV with two other model-induced annotation methods to showcase the effectiveness and efficiency of our proposed approach. In Section 6.2, we validate the data quality constructed via AutoCV as described in Section 5.2.
6.1 Different Process Labelling Strategy
Aside from the labeling method defined by Eq. (3) in AutoCV, another labeling strategy is the Monte Carlo sampling estimation (MCS), as described in  [17, 18]. To better demonstrate the effect of our method, we make a comparison with the approaches proposed by [17, 18], and conduct experiments on mathematical benchmarks. The results are shown in Table 8. We follow the experimental settings described in their work to ensure a fair comparison. More implementation details are provided in Section F.3.
Table 8: Comparison of process labeling methods’ performance across different response generators on GSM8K and MATH datasets. The table evaluates the Pass@5, Self-Consistency (Self-Cons.), and response selection performance of models fine-tuned using process annotations labeled by MCS and our proposed method (AutoCV).
@@ 6.1 Different Process Labelling Strategy
Pass@5 Self-Cons. Process (MCS) Process (AutoCV) Pass@5 Self-Cons. Process (MCS) Process (AutoCV)
@@
Table 9: Comparison of annotation costs between MCS and AutoCV for process labeling on the GSM8K and MATH datasets. Annotation cost represents the processed tokens into a model when generating process annotations, encompassing both input and output tokens.
@@
#Steps(Avg.) #Steps(Overall) #Tokens(Avg.) #Tokens(Overall) Process (MCS) Process (AutoCV)
@@
The experimental results shown in Table 8 suggest that our proposed method for process labeling, which relies on detecting changes in model confidence, performs competitively with the MCS method from [18, 17]. In some cases, our method even outperforms the MCS method, especially on the more challenging MATH benchmark.
It is important to note that, as mentioned in the analysis and shown in Table 9, our method is computationally more efficient than the MCS method, as it does not require generating costly multiple samples to label each reasoning step. This computational efficiency advantage of our method could be particularly beneficial for large-scale applications or scenarios with limited computational resources.
In summary, the experimental results and analysis demonstrate that our proposed process labeling method offers a promising alternative to the MCS method, providing competitive or better performance while being computationally more efficient. Moreover, further research can benefit from the advantages of both approaches, potentially achieving better overall performance and efficiency.
@@
The experimental results in Table 10 reveal that retraining the model with process supervision from AutoCV still yields better performance than self-consistency across three different response generators. We also noticed a small performance gap between the PSV and OSV. It is worth noting that our PSV was trained using data from the OSV. The small performance gap between the PSV and OSV models demonstrates that the relabeled process-supervised training method successfully inherits information from the outcome-supervised model without requiring ground truth annotations. This ablation study further provides quality assurance for automatic process labeling via AutoCV.
@@ 7 Conclusion
In this paper, we propose a novel method for automated process labeling via confidence variation (AutoCV) to enhance the reasoning capabilities of LLMs by detecting relative changes in model confidence. It combines the strengths of output supervision and process supervision to annotate reasoning steps automatically. Extensive experiments demonstrate that AutoCV significantly enhances the precision and scalability of the verifier models in various reasoning tasks, ranging from mathematical to commonsense reasoning. AutoCV could considerably enhance existing LLMs’ performance while drastically reducing the need for intensive computation and manual intervention.
For future work, we aim to utilize the automatically constructed PSV to supervise the generator using step-wise proximal policy optimization, to enhance the accuracy of the generator’s output during greedy decoding without the need for subsequent reranking. This avenue of research could lead to even more advancements in the capabilities of LLMs and their application in reasoning tasks. The limitations and broader impact of the paper are discussed in Appendix A and B.
@@ Appendix A Limitations
AutoCV is a promising solution for enhancing the reasoning capabilities of LLMs. However, it is important to acknowledge several potential limitations of the method. Firstly, while AutoCV aims to reduce the need for manual intervention, there is still a risk of inaccurate annotations. The relative confidence variation used to produce process annotations is an estimation and may not always accurately represent the actual correctness of a reasoning step. This could compromise the quality of the annotations and, in turn, the effectiveness of the method. Secondly, the success of AutoCV is heavily dependent on the performance of the verifier. If the model is not accurate enough in its step-level scores, the quality of the process annotations generated by AutoCV could be compromised. Thirdly, AutoCV is specifically designed to improve the reasoning capabilities of LLMs. Therefore, its applicability may be limited to tasks that involve complex multi-step reasoning. It is unclear how well the method will scale or generalize to other tasks or domains that do not involve intensive reasoning. This is an important consideration for future research and development of the method.
@@ Appendix B Broader Impact
The proposed AutoCV has the potential to bring about several positive societal impacts. By enhancing the reasoning capabilities of LLMs, AutoCV can lead to more accurate and reliable information, which in turn can support better decision-making in various sectors, including healthcare, education, and finance. Moreover, AutoCV combines the strengths of output supervision and process supervision to automatically annotate reasoning steps, significantly reducing the time, effort, and cost associated with manual annotation. This makes the process of training LLMs more efficient and accessible. Additionally, the process supervision data generated by AutoCV can improve the performance and scalability of verification models, allowing for the development of more complex and sophisticated LLMs capable of handling a wider range of tasks and applications.
@@
However, AutoCV also presents several potential negative societal impacts. The automation of the annotation process could lead to job displacement for individuals currently employed in this role. There is also a risk that AutoCV and the enhanced LLMs could be misused, for instance, to spread misinformation or manipulate public opinion. The increased reliance on LLMs for decision-making could potentially result in a decrease in critical thinking and problem-solving skills among individuals. Furthermore, the use of LLMs in various sectors could lead to privacy and security issues, as these models often require large amounts of data for training.

### v4

@@ 4.2.2 Quantitative Results
We introduce three metrics for a thorough evaluation of math calculation error detection: Precision (Prec.), which calculates the proportion of samples with correct final answers but exhibiting hallucinatory errors during the reasoning process; Recall, which determines the proportion of samples with math calculation errors that the OSV model successfully identifies through step-level confidence changes; F1-score, which gauges the verifier’s overall efficacy. In Table 5, we explore how different threshold ($\theta$) values affect the precision, recall, and F1-score for math calculation error detection.
Table 5: Process Calculation Error Detection Performance with Varying Threshold ($\theta$) Values.
@@
The results in Table 5 demonstrate that our method using step-level confidence change effectively detects calculation errors across threshold values from - 0.5 to - 0.9. As the threshold becomes more negative (stricter for labeling errors), the precision increases, indicating higher precision in identifying true errors. However, the recall decreases, meaning fewer actual errors are caught. Importantly, the F1-score, balancing precision and recall, remains relatively stable across thresholds. This demonstrates that our method strikes a good balance between detecting real errors and minimizing incorrect flagging of valid calculations. Overall, our detection method is effective and robust, performing well over a range of thresholds without significantly compromising overall detection quality.
We note that setting $\theta$ = - 0.5 in our detection methods helps maintain a balance between precision and recall, which can ensure a balanced distribution of labeled “incorrect” and “correct” responses.
Validation and Foundation for AutoPSV
Our empirical validation of the OSV model encompasses two key aspects: its efficacy in response selection (Section 4.1) and its capability in detecting calculation errors (Section 4.2). These experimental results provide a robust foundation for automating process annotations using AutoPSV. Furthermore, Theorem 1 establishes the theoretical framework for utilizing OSV to estimate the probability of reaching correct final answers from any given intermediate reasoning step. This convergence of theoretical guarantees and empirical evidence provides the methodological groundwork for applying AutoPSV to generate process-supervised training data in our subsequent experiments.
@@ 5.1 Experimental Setup
Models:
We selected three instruction-tuned LLMs of varying sizes, ranging from 7 billion to over 70 billion parameters, to serve as the response generator. Specifically, we used Mistral-Instruct-7B (Mistral-Instruct), Mixtral-8x7B-Instruct-v0.1 (Mixtral-Instruct), and Qwen-72B-Chat (Qwen).
Datasets:
Our evaluation encompasses five benchmarks across two reasoning domains: For mathematical reasoning, we include GSM8K [21], a dataset containing math word problems requiring multi-step reasoning, and MATH [36], composed of high school-level competition problems covering a range of math subjects. For commonsense reasoning, we use HellaSwag [37], a dataset for physically situated commonsense reasoning, Winogrande [38], fill-in-the-blank problems requiring commonsense pronoun resolution and ANLI [39], a dataset for natural language understanding and reasoning.
Evaluation:
For evaluation, we follow the methodology outlined in [40] to ensure consistency across benchmarks. Our protocol involves:
(i) Our object is to select the correct answer from five candidate responses.
(ii) For OSV models, we use the final value assigned to the whole solution for evaluation, while for PSV models, we utilize the product of step-level scores as the aggregation function.
(iii) To obtain more reliable pass@k results, we implement the estimation method described in [41]. This involves generating n samples per task (where n > k) and assessing the number of correct samples that pass unit tests. We then calculate the unbiased estimator for pass@k. For self-consistency (Self-Cons.) and verifier results, we randomly select k out of n samples and perform separate calculations. All results are reported with an accuracy of ±0.1 at a 95% confidence level.
Additional details regarding generation hyperparameters and different aggregation functions are provided in Section F.2.
@@ 5.2 Enhanced LLMs Reasoning via Process Supervision
To evaluate the efficacy and scalability of our proposed approach (detailed in Section 5.2), we conducted comprehensive experiments across mathematics and commonsense reasoning tasks using five diverse datasets.
Our experimental framework employs an autonomous process-supervision data annotation method where we calculate $\Delta_{conf}^{t}$ based on model confidence from OSV, using a threshold of $\theta$ = -0.5. This annotated data is then utilized to continuously fine-tune the OSV model, resulting in our enhanced OSV + PSV model.
We evaluate three distinct approaches: (i) Self-Consistency (Self-Cons.): Our baseline approach (ii) Outcome-supervised verifier (OSV): Our initial verification model (iii) Process-supervised enhanced verifier (OSV + PSV): Our proposed enhancement. For each approach, we assess Pass@5 performance, which represents the upper limit of achievable performance on these benchmarks.
Mathematics Reasoning:
As shown in Table 6, the process-supervised enhanced verifier demonstrates superior performance over the outcome-supervised verifier and Self-Consistency models for all evaluated response generators on GSM8K. For the MATH benchmark, the process-supervised enhanced verifier outperforms the other two approaches for Mistral-Instruct and Mixtral-Instruct, but it is slightly less effective than the Self-Consistency model when applied to Qwen-72b.
@@
Mistral-Instruct 69.90 50.03 61.18 61.41 7.7 1.64 5.10 5.30
Mixtral-Instruct 82.30 69.06 74.91 76.04 22.80 10.66 15.2 16.92
Qwen 91.13 81.27 84.91 85.15 56.10 40.10 38.94 39.36
Commonsense Reasoning:
According to Table 7, OSV + PSV again leads to the best results among the three methods for each response generator tested on HellaSwag. For Winogrande, Mistral-Instruct paired with OSV + PSV achieves the highest performance, whereas, for Mixtral-Instruct and Qwen-72b, the original OSV without process supervision has a marginal advantage. When looking at the results of the ANLI benchmark, OSV + PSV is the highest-performing method for Mistral-Instruct and Mixtral-Instruct. Despite this, for Qwen-72b, the OSV model alone falls slightly behind the integrated OSV + PSV.
@@
Mistral-Instruct 76.84 40.30 73.81 74.45 91.16 58.64 79.16 79.98 73.4 45.6 59.8 59.3
Mixtral-Instruct 84.05 73.67 82.83 83.62 79.16 68.75 73.40 73.88 68.4 59.0 62.9 64.0
@@
Conclusion:
Our experimental results indicate that the process-supervised enhanced verifier (OSV + PSV) consistently outperforms or matches the baseline models across most mathematical and commonsense reasoning tasks. By leveraging automatic process annotations, our approach enhances the model’s capacity to verify reasoning processes, resulting in improved accuracy and robustness across a wide range of benchmarks and response generators.
@@ 6 Analysis
In Section 6.1, we compare our process annotation method, AutoPSV with two other model-induced annotation methods to showcase the effectiveness and efficiency of our proposed approach. In Section 6.2, we validate the data quality constructed via AutoPSV as described in Section 5.2.
6.1 Advantages of AutoPSV in Labeled and Unlabeled Settings
Aside from the labeling method defined by Eq. (3) in AutoPSV, another labeling strategy is the Monte Carlo Tree sampling estimation (MCTS), as described in  [17, 18]. To better demonstrate the effect of our method, we make a comparison with this approach and conduct experiments on mathematical benchmarks across both labeled and unlabeled settings (i.e., whether the ground-truth answers to the questions are provided).
Performance in Labeled Settings
We first evaluate the performance of AutoPSV in labeled settings. We follow the experimental settings described in [17, 18] to ensure a fair comparison. More implementation details are provided in Section F.3. Table 8 presents a comparison of process labeling methods across different response generators on the GSM8K and MATH datasets.
Table 8: Comparison of process labeling methods’ performance across different response generators on GSM8K and MATH datasets. The table evaluates the Pass@5, Self-Consistency (Self-Cons.), and response selection performance of models fine-tuned using process annotations labeled by MCTS and AutoPSV
@@ 6.1 Advantages of AutoPSV in Labeled and Unlabeled Settings
Pass@5 Self-Cons. Process (MCTS) Process (AutoPSV) Pass@5 Self-Cons. Process (MCTS) Process (AutoPSV)
@@
The experimental results shown in Table 8 suggest that our proposed method for process labeling, which relies on detecting changes in model confidence, performs competitively with the MCTS method from [18, 17]. In some cases, our method even outperforms the MCTS method, especially on the more challenging MATH benchmark.
A key advantage of AutoPSV is its computational efficiency. As shown in Table 9, our method requires significantly fewer tokens for process labeling compared to MCTS-based methods.
Table 9: Comparison of annotation costs between MCTS and AutoPSV for process labeling on the GSM8K and MATH datasets. Annotation cost represents the processed tokens into a model when generating process annotations, encompassing both input and output tokens.
@@
#Steps(Avg.) #Steps(Overall) #Tokens(Avg.) #Tokens(Overall) Process (MCTS) Process (AutoPSV)
@@
This efficiency stems from AutoPSV’s ability to generate process annotations without requiring multiple samples for each reasoning step, making it particularly suitable for large-scale applications or scenarios with limited computational resources.
Performance in Unlabeled Settings
To further demonstrate the flexibility of AutoPSV, we evaluate its performance in unlabeled settings. We generated an additional dataset of 7,000 unlabeled math problems using the Evol-Instruct method from WizardLM [42]. These problems, created by LLMs without accompanying gold solutions, represent a challenging scenario for traditional supervision methods. We then conducted an experiment incorporating these unlabeled questions alongside the GSM8K dataset. Table 10 compares various methods across different response generators. In this context, OSV+PSV (GSM8K) refers to the original AutoPSV setting, while OSV+PSV (GSM8K+WizardLM) includes process annotations sourced from both GSM8K and WizardLM unlabeled questions. Notably, both MCTS and OSV-only training cannot leverage these unlabeled data, highlighting another key advantage of AutoPSV.
Table 10: Performance enhancement of our proposed AutoPSV method in unlabeled settings, where both MCTS and OSV-only training are unable to utilize unlabeled data.
Response Generator Pass@5 Self-Cons. OSV (GSM8K) MCTS (GSM8K) OSV+PSV (GSM8K) OSV+PSV (GSM8K+WizardLM)
Mistral-Instruct 69.90 50.03 61.18 60.82 61.41 63.11
Mixtral-Instruct 82.30 69.06 74.91 75.10 76.04 78.15
Qwen 91.13 81.27 84.91 84.85 85.15 86.77
The results demonstrate that the addition of unlabeled data leads to noticeable improvements across all response generators. For instance, the performance of Mistral-Instruct improves from 61.18 (OSV) to 63.11 (OSV+PSV with GSM8K+WizardLM). These results further underscore the value of the AutoPSV approach, particularly its ability to effectively utilize unlabeled data for enhanced performance.
In summary, AutoPSV offers several key advantages over MCTS-based methods:
Consistent Performance with Computational Efficiency: AutoPSV demonstrates robust, consistent improvements across different response generators and datasets. Its ability to efficiently generate process annotations—without the need for extensive sampling like MCTS—makes it particularly well-suited for large-scale applications and environments with limited computational resources.
Leveraging Unlabeled Data for Enhanced Performance: Unlike MCTS, which relies on ground truth labels, AutoPSV can effectively utilize unlabeled data. This capability not only enhances model performance in real-world settings but also offers scalability in scenarios where labeled data is scarce. The flexibility to harness unlabeled data ensures that AutoPSV can drive significant improvements even in data-constrained situations.
@@
The experimental results in Table 11 reveal that retraining the model with process supervision from AutoPSV still yields better performance than self-consistency across three different response generators. We also noticed a small performance gap between the PSV and OSV. It is worth noting that our PSV was trained using data from the OSV. The small performance gap between the PSV and OSV models demonstrates that the relabeled process-supervised training method successfully inherits information from the outcome-supervised model without requiring ground truth annotations. This ablation study further provides quality assurance for automatic process labeling via AutoPSV. Moreover, OSV training is limited to labeled datasets, while AutoPSV demonstrates superior performance by utilizing both labeled and unlabeled data as shown in Table 10. This comparison further highlights the versatility and effectiveness of AutoPSV in real-world scenarios where ground truth annotations may be scarce or unavailable.
@@ 7 Conclusion
In conclusion, we propose a novel method for automatic process labeling in LLMs by detecting relative changes in model confidence. Our experimental results demonstrate that AutoPSV significantly enhances the precision and scalability of the verifier models in various reasoning tasks, ranging from mathematical to commonsense reasoning. AutoPSV therefore has the potential to considerably enhance existing LLMs’ performance while drastically reducing the need for intensive computation and manual intervention. For future work, we aim to utilize the automatically constructed PSV to supervise the generator using step-wise proximal policy optimization, to enhance the accuracy of the generator’s output during greedy decoding without the need for subsequent reranking. This avenue of research could lead to even more advancements in the capabilities of LLMs and their application in reasoning tasks. The limitations and broader impact of the paper are discussed in Appendix A and B.
@@ Appendix A Limitations
AutoPSV is a promising solution for enhancing the reasoning capabilities of LLMs. However, it is important to acknowledge several potential limitations of the method. Firstly, while AutoPSV aims to reduce the need for manual intervention, there is still a risk of inaccurate annotations. The relative step-level confidence change used to produce process annotations is an estimation and may not always accurately represent the actual correctness of a reasoning step. This could compromise the quality of the annotations and, in turn, the effectiveness of the method. Secondly, the success of AutoPSV is heavily dependent on the performance of the verifier. If the model is not accurate enough in its step-level scores, the quality of the process annotations generated by AutoPSV could be compromised. Thirdly, AutoPSV is specifically designed to improve the reasoning capabilities of LLMs. Therefore, its applicability may be limited to tasks that involve complex multi-step reasoning. It is unclear how well the method will scale or generalize to other tasks or domains that do not involve intensive reasoning. This is an important consideration for future research and development of the method.
@@ Appendix B Broader Impact
The proposed AutoPSV has the potential to bring about several positive societal impacts. By enhancing the reasoning capabilities of LLMs, AutoPSV can lead to more accurate and reliable information, which in turn can support better decision-making in various sectors, including healthcare, education, and finance. Moreover, AutoPSV combines the strengths of output supervision and process supervision to automatically annotate reasoning steps, significantly reducing the time, effort, and cost associated with manual annotation. This makes the process of training LLMs more efficient and accessible. Additionally, the process supervision data generated by AutoPSV can improve the performance and scalability of verification models, allowing for the development of more complex and sophisticated LLMs capable of handling a wider range of tasks and applications.
@@
However, AutoPSV also presents several potential negative societal impacts. The automation of the annotation process could lead to job displacement for individuals currently employed in this role. There is also a risk that AutoPSV and the enhanced LLMs could be misused, for instance, to spread misinformation or manipulate public opinion. The increased reliance on LLMs for decision-making could potentially result in a decrease in critical thinking and problem-solving skills among individuals. Furthermore, the use of LLMs in various sectors could lead to privacy and security issues, as these models often require large amounts of data for training.